OpenStudy (anonymous):

x(10x-2)=3 use quadratic formula to solve this.

OpenStudy (anonymous):

Please help me with this someone? Please, i need to get this done in four hours.

OpenStudy (across):

Expand the expression and rearrange\[10x^2-2x-3=0\]

OpenStudy (across):

I suppose you already know how to complete the square? Seems the purpose of this homework is to train you on it.

OpenStudy (anonymous):

i have what you typed so far

OpenStudy (anonymous):

she didnt teach us to complete the square we skipped that

OpenStudy (anonymous):

she said we have to solve using quadratic formula

OpenStudy (across):

Oh! Okay.

OpenStudy (across):

Well, you can see that a = 10, b=-2 and c = -3.

OpenStudy (anonymous):

yes

OpenStudy (across):

\[\frac{-(-2)\pm\sqrt{(-2)^2-4(10)(-3)}}{2(10)}\]

OpenStudy (across):

\[\frac{2\pm\sqrt{4+120}}{20}\]

OpenStudy (across):

\[\frac{1}{10}\pm\frac{\sqrt{31}}{10}\]

OpenStudy (anonymous):

How did you get that

OpenStudy (across):

\[\frac{2}{20}=\frac{1}{10}\]and

OpenStudy (across):

\[\frac{\sqrt{124}}{20}=\frac{\sqrt{4\cdot31}}{20}=\frac{2\sqrt{31}}{20}=\frac{\sqrt{31}}{10}\]

OpenStudy (anonymous):

\[2+-2\sqrt{31}/10\]

OpenStudy (anonymous):

This part bugs me a lot

OpenStudy (anonymous):

/20 i mean

OpenStudy (across):

\[\frac{2\sqrt{31}}{20}=\frac{2}{20}\sqrt{31}=\frac{1}{10}\sqrt{31}=\frac{\sqrt{31}}{10}\]

OpenStudy (anonymous):

what happened to the other 2?

OpenStudy (asnaseer):